a) Find the linear and quadratic approximations of sin(0.96Ï€) tan(0.26Ï€) + (0.96)2(0.26)

b) Compare these approximations with the actual value.


1
Expert's answer
2021-09-23T14:15:14-0400

"f(x,y)=sin(\pi x)tan(\pi y)+x^2y"

"x_0=0.96,\ y_0=0.26"


a)

Linear approximation:

"L(x,y)=f(x_0,y_0)+f_x(x_0,y_0)(x-x_0)+f_y(x_0,y_0)(y-y_0)"

"f_x(x,y)=\pi cos(\pi x)tan(\pi y)+2xy"

"f_y(x,y)=\frac{\pi sin(\pi x)}{cos^2(\pi y)} +x^2"

"f(x_0,y_0)=sin(0.96\pi )tan(0.26\pi )+0.96^2\cdot 0.26=0.37"

"f_x(x_0,y_0)=\pi cos(0.96\pi )tan(0.26\pi )+2\cdot 0.96\cdot 0.26=-2.82"

"f_y(x_0,y_0)=\frac{\pi sin(0.96\pi )}{cos^2(0.26\pi )} +0.96^2=1.77"

"L(x,y)=0.37-2.82\cdot(x-0.96)+1.77\cdot(y-0.26)"

"L(x,y)=2.62-2.82x+1.77y"


Quadratic approximation:

"Q(x,y)=f(x_0,y_0)+f_x(x_0,y_0)(x-x_0)+f_y(x_0,y_0)(y-y_0)+"

"+\frac{1}{2}f_{xx}(x_0,y_0)(x-x_0)^2+f_{xy}(x_0,y_0)(x-x_0)(y-y_0)+"

"+\frac{1}{2}f_{yy}(x_0,y_0)(y-y_0)^2"


"f_{xx}(x,y)=-\pi^2 sin(\pi x)tan(\pi y)+2y"

"f_{yy}(x,y)=\frac{2\pi^2 sin(\pi x)sin(\pi y)}{cos^3(\pi y)}"

"f_{xy}(x,y)=\frac{\pi^2 cos(\pi x)}{cos^2(\pi y)} +2x"